A truck enters a highway driving 60 mph. A car enters the highway at the same place 13 minutes later and drives 66 mph in same direction. From the time the car enters the highway, how long will it take the car to pass the truck
Found 2 solutions by richwmiller, Theo:
Answer by richwmiller(17219)

Gaining speed 66 -60= 6 mph
Headstart 60*0.21666667 =13.0 miles
Time to catch up 13.0/6=2 1/6 hours= 2 hours 10 minutes
Algebraic solution
60(t+0.21666667)=66t
60t+13.0=66t
13.0=66t-60t
13.0=6t
t=13.0/6=2 1/6 hours= 2 hours 10 minutes
Distance to meeting 2.16666667*66=143 miles
code drt1

Answer by Theo(13342)

rate * time = distance

13 minutes divided by 60 = .2166... hours.

rate * time = distance for the truck is 60 * x = d, where d is the distance the car travels and x is the time it takes to travel that far in hours.

rate * time = distance for the car is 66 * (x-.2166...) = d, where d is the distance the car travels and (x-.2166...) is the time it takes to travel that far in hours.

when the car has caught up with the truck, they have both traveled the same distance.

since rate * time = distance, this allows you to set rate * time for the truck equal to rate * time for the car.

you get:

60 * x = 66 * (x-.2166...)

solve for x in equation to get x = 2.3833... hours.

the truck has traveled for 2.3833... hours and the car has traveled for 2.3833... - .2166... hours which is equal to 2.166... hours.

since rate * time = distance, then:

the distance the truck has traveled is 60 * 2.3833.... = 143 miles.

the distance the car has traveled is 66 * 2.166... = 143 miles.

It will take the car 2.166... hours to pass the truck and the car will do it in 143 miles.

2.166... hours is equivalent to 2.166... * 60 minutes which is equal to 130 minutes.